OpenStudy (anonymous):

HELP PLEASE Find the zeros of f(x)=(x+5)^5(x-3)^6 and state the multiplicity.

OpenStudy (campbell_st):

to find the zeros you need to solve x + 5 = 0 and x - 3 = 0 the multiplicity of each zero will be the power of the factor (x + 5)^5 has multiplicity of 5... I'll let you work out the multiplicity of the other factor. hope it helps

OpenStudy (mathstudent55):

Set the function equal to zero: (x + 5)^5(x - 3)^6 = 0 Because of the exponent of 5 on x + 5, you have 5 of those factors: (x + 5)(x + 5)... 5 equal factors Because of the exponent of 6 on x - 3, you have 6 of those factors (x - 3)(x - 3) ... 6 equal factors Solve (x + 5)(x - 3) = 0 x + 5 = 0 or x - 3 = 0 x = -5 or x = 3 Because the x + 5 has a 5 exponent, the multiplicity is 5 Because the x - 3 has an exponent of 6, the multiplicity is 6 So the answer is x = -5 multiplicity 5 or x = 3 multiplicity 6

OpenStudy (anonymous):

the other has a multiplicity of 6? so answer would be???

OpenStudy (anonymous):

THANKS!

OpenStudy (campbell_st):

thats correct... you just need to find the roots... by solving the equations... x + 5 = 0 and x - 3 = 0